OpenStudy (anonymous):

y = sinx cosx tanx^3/ √x, find dy/dx with logarithmic differentiation

OpenStudy (agreene):

\[y=\frac{\sin(x)\cos(x)\tan(x^3)}{\sqrt{x}}\] is that the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take the log, get \[\ln(\sin(x))+\ln(\cos(x))+\ln(\tan(x))-\frac{1}{2}\ln(x)\] then take the derivative piece by piece then multiply by the original function

OpenStudy (anonymous):

\[\ln(\sin(x))+\ln(\cos(x))+\ln(\tan(x^3))-\frac{1}{2}\ln(x)\] forgot the x cubed! sorry